Operation Summary for November 18, 2009 - November 20, 2009 |
Wednesday Nov. 18 The day shift began with the Tevatron A-E enclosures prepared for access, with Operations monitoring the Recycler stash (264.7E10) and the Antiproton Source stack (12.3E10), with the Neutron Therapy Facility treating patients, but with the rest of Linac down for RF station (LRF5) repairs, with Booster in access, and with the Main Injector (MI) and Recycler operating normally, with the NuMI and MiniBooNE experiments off due to Linac and Booster work, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments in access. During the day shift, the TeV was in access all morning for the sector A2 cryo system lead repair and the A1 dipole magnet alignment, among other things. The A-E sectors access lasted until 1 PM with the cooldown of A1 house lasting through to the end of the shift. During the F-Sector access there was no beam to users, so Pbar tunnels were accessed for work as well. Beam was re-established to Pbar, NuMI, and MiniBooNE users at 1:30 PM. MI RF problems held off NuMI beam from 2:45 until 3:15 PM. During the evening shift, the A1 cryo system cooldown was completed by 6 PM. Operations conducted TeV ramping and orbit checks until 11 PM. A controls crate power supply failure at F0 held off the TeV and halted the HEP shot setup. Operations called in cryo personnel to replace the failed supply. Pbar stacked and the Recycler stashed all shift. Thursday Nov. 19 The midnight shift began with TeV in a wet squeeze, with Operations monitoring the stash (418.7E10), with Pbar stacking (13.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, cryo personnel had the TeV sector F0 control crate power supply replaced by 1:15 AM. Operations put the TeV into shot setup. Operations established store 7370 at 3 AM with an average initial luminosity (AIL) of 272E30. Pbar stacked and the Recycler stashed all shift. During the day shift, at 10:54 AM, the H- Source extractor began to fail. Preaccelerator personnel accessed and determined that the source racetrack needed to be replaced. This work continued through to the end of the shift. Instrumentation personnel accessed the Pbar vault at AP0 to replace a target SEM. Booster personnel began work on their extraction kickers. TeV personnel adjusted the D0 interaction point. During the evening shift, Preaccelerator personnel had the racetrack replaced by 5:50 PM, but a problem with the cesium boiler required another access; they found and replaced a bad power supply by 11 PM and then began the source recovery process. Friday Nov. 20 The midnight shift began with Operations monitoring store 7370 (50.77E30), the stash (280.2E10), and the Pbar stack (7.2E10), and with NuMI and MiniBooNE off due to the H- Source recovery. During the midnight shift, Operations reported that the H- Source had warmed up and stabilized by 5 AM and then turn-on and tuning continued through to the end of shift. The Plans for Friday and the Weekend The plans for today are to bring the H- Source online and then resume stacking, stashing, and storing. The Run Coordinator will allow Linac, Booster, MI and Recycler tuning and studies. There is a MI vacuum problem that might require an access to fix. Switchyard personnel will send beam to MTest on Monday and Tuesday. |
|
More Information Comments and Suggestions |
| last modified 11/20/2009 email Fermilab |
| Security, Privacy, Legal | |